Your friend owns a small business and she asks for your advice. For the past couple of​ years, her company has extended credit to its customers. She wonders how well her company manages its accounts receivable. During the most recent​ year, your​ friend's company had net credit sales of $743,000. Net Accounts receivable at the beginning of the year was $82,000. Ending net Accounts receivable was $77,000.

The​ company's credit terms are net 30.

What should you tell your friend regarding how well accounts receivable is​ managed?

Begin by calculating the​ business's accounts receivable turnover ratio for the year. ​(Round the accounts receivable turnover ratio to two decimal​ places, X.XX.)

Net credit sales / Average net accounts receivable = Accounts receivable turnover ratio

$743,000 / $79,500 = 9.35

With credit terms of 30​ days, you would expect to have an accounts receivable turnover of approximately (5 ,8, 12, 15, 20). When comparing this number with the accounts receivable turnover ratio calculated in the preceding​ step, you should tell your friend that it appears that the company is not managing its accounts receivable very well.
Answer #1

Net credit sales = $743,000. Net Accounts receivable at the beginning of the year was $82,000. Ending net Accounts receivable was $77,000.

Average accounts receivable = (Beginning accounts receivable + Ending accounts receivable)/2

= (82,000 + 77,000)/2

= $79,500

Accounts receivable turnover ratio =  Net credit sales / Average net accounts receivable

= $743,000 / $79,500

Accounts receivable turnover ratio = 9.35

Credit period allowed by the company = 30 days

Average collection period = 365/Accounts receivable turnover ratio

30 = 365/Accounts receivable turnover ratio

Accounts receivable turnover ratio = 365/30

= 12

With credit terms of 30​ days, expected accounts receivable turnover of approximately = 12

When accounts receivable turnover ratio is 9.35, average collection period will be calculated as under:

Average collection period = 365/Accounts receivable turnover ratio

= 365/9.35

= 39 days

If Company's Accounts receivable turnover ratio is 9.35, company is collecting cash from customers in 39 days while the credit period allowed by it, is 30 days.

Since average collection period (39 days) is more than credit period allowed (30 days), hence company is not managing its accounts receivable very well.

Kindly comment if you have any doubt. Thanks.